OpenStudy (anonymous):

Divide. (23z^3v^7-27z^7v^7) / (-3z^4v^3) Simplify your answer as much as possible.

OpenStudy (anonymous):

@Dankknight @DoShKa_SyRiA

OpenStudy (anonymous):

@jim_thompson5910 @Hero @zepdrix @abb0t

OpenStudy (anonymous):

help

OpenStudy (abb0t):

Can you rewrite it using the equation editor under this little box? That way, it is clear for both the helper and the helpee.

OpenStudy (anonymous):

\[(23z^3v^7-27z^7v^7)\div(-3z^4v^3)\]

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@ThatAPLyfe

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

use this quotient property of powers \(a^m/a^n=a^{m-n}\)

OpenStudy (xapproachesinfinity):

first thing is you can separate the fraction into two fractions using arithmetic rules of fractions \(\Large \frac{23z^3v^7-27z^7v^7}{-3z^4v^3}=\frac{23z^3v^7}{-3z^4v^3}-\frac{27z^7v^7}{-3z^4v^3}\)

OpenStudy (xapproachesinfinity):

i used this property \(\huge\frac{a+b}{c}=\frac{a}{c}+\frac{b}{c}\)